OpenStudy (jennychan12):

Find integral from 0 to 1/sqrt2 of arccosx/(1+x^2)^1/2.

OpenStudy (jennychan12):

\[\int\limits_{0}^{1/\sqrt{2}} \frac{ arc \cos x }{ \sqrt{1-x^2} }dx\]

OpenStudy (jennychan12):

this is the work they have. but what i don't get is how they got the 1/2

OpenStudy (anonymous):

are u there this is simple to solve...

OpenStudy (anonymous):

when u see the derivative of the function in your integral then u just use the simple porwer rule

OpenStudy (anonymous):

so in ur case u have the derivative of arccosx in the denominator, hence, all u have to do is use 1/2 arcsin^2(x). recall the basic 1/n+1x^(n+1)

OpenStudy (anonymous):

also notice the arc cosx became arc sinx

OpenStudy (anonymous):

so u can use the u sub

OpenStudy (anonymous):

one of the ways to solve this problem or refer to the table

OpenStudy (anonymous):

the 1/2 should be there try the u sub. u'll see u'll get 1/2 u^2

OpenStudy (jennychan12):

oh wait. i think i see what your saying @mathsmind cuz it's basically just integral u du right?

OpenStudy (anonymous):

i will show u in details

OpenStudy (jennychan12):

no, no, i understand thanks :)

OpenStudy (anonymous):

oh ok yes u got what i said

OpenStudy (anonymous):

so keep that in mind and remeber it like ur name

OpenStudy (jennychan12):

-_- what?

OpenStudy (anonymous):

it is very important to search for the derivative in an integral before u solve it

OpenStudy (anonymous):

i mean remeber the concept that we just said like ur name

OpenStudy (anonymous):

that will help u solve integrals in ur mind

OpenStudy (anonymous):

i never saw the answer u posted because i don't open image files

OpenStudy (jennychan12):

what do integrals have to do with my name?

OpenStudy (anonymous):

this is a metaphor i meant that remeber this rule like u know ur name, the same way if someone ask u about ur name, for example if one ask u what is ur name, you know the answer straight way, so know this rule in the same manner...

OpenStudy (jennychan12):

ohhh ok. sorry. i'm slow thanks for your tip :)